Grade 12th passMechanics

A wheel with a fixed angle rotating so that the anstantaneous velocity of a referance line painted along a radious is w=At+Bt2 when A=6.2 rad/s2 and B=8.7 rad/s2.
  1. If the referance line is instantly at ɵ=0 when t=0 finds its angular position when t=2.0 sec.
  2. What is the instantaneous angular accleration of the reference line at t=0.50s

Hridoy
8 Years agoGrade 12th pass

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
w = d\Theta/dt = At + Bt2
Let angular position at t = 2sec is \Theta then w is geven by
(\Theta-0)/(2-0) = 6.2*2 + 8.7*(2)2
\Theta/2 = 47.2
So \Theta = 94.4 rad.
angular acceration will be given by
dw/dt = A + 2Bt
So its value at 0.5 sec will be
6.2 + 2*8.7*0.5 = 14.9 rad/s2.
Hope it clears.
If you like answer then please approve it.